# CANDLE: Iterative Conceptualization and Instantiation Distillation from Large Language Models for Commonsense Reasoning

### v1

@@ Acknowledgements
The authors of this paper were supported by the NSFC Fund (U20B2053) from the NSFC of China, the RIF (R6020-19 and R6021-20), and the GRF (16211520 and 16205322) from RGC of Hong Kong. We also thank the support from the UGC Research Matching Grants (RMGS20EG01-D, RMGS20CR11, RMGS20CR12, RMGS20EG19, RMGS20EG21, RMGS23CR05, RMGS23EG08).
@@ Appendix A Distillation Details
These prompts are consistent with our descriptions in Section 4.1 and Section 4.2, where the task description is first presented, followed by human-authored examples, and finally, the event we want to conceptualize or instantiate.
Additionally, we introduce some generation settings when prompting LLMs. For ChatGPT, we access it through the official OpenAI APIs111https://chat.openai.com/. The code of the accessed version is gpt-3.5-turbo-0613. We set the temperature to 0.7 and the maximum length for generated tokens to 200. To conceptualize all events in ATOMIC into 20 conceptualizations each, the time required for the distillation process is approximately ten days.
For LLAMA2, we access it via the Huggingface222https://huggingface.co/meta-llama/Llama-2-13b-chat-hf Library. The code of the accessed model is meta-llama/Llama-2-13b-chat-hf. When prompting, we use the Top-k sampling decoding strategy and set . We set the maximum length of generated tokens to 200. The models are hosted on sixteen NVIDIA-V100 GPUs, and the time required to distill the entire dataset is approximately one month.
@@
Furthermore, we employ BLEU soft uniqueness Zhu et al. (2018); West et al. (2022) to evaluate the distinctiveness of the newly introduced head events resulting from our chain of conceptualization and instantiation. To determine uniqueness, we define an instantiated head event, referred to as , as unique if , where represents the original head event in ATOMIC. The threshold of 0.5 is an empirical threshold. Our empirical results demonstrate that 78.6% of the instantiated events are unique compared to ATOMIC, highlighting the effectiveness of CANDLE in enhancing the semantic coverage of the CSKB.
Future works can experiment with applying CANDLE distillation framework to other downstream tasks, such as complex-query answering Bai et al. (2023), script generation Yuan et al. (2023), pragmatics Ou et al. (2023a, b), analogical reasoning Cheng et al. (2023), and E-Commerce understanding Yu et al. (2023).
Appendix B Dataset Descriptions
This section covers additional details and statistics of datasets and benchmarks used in downstream task evaluations. Please refer to the corresponding subsection for each downstream task.
@@ Appendix B Dataset Descriptions
B.1 CSKB Conceptualization
In CSKB Conceptualization tasks, we use the AbstractATOMIC He et al. (2022a) dataset as the evaluation benchmark. It is a benchmark dataset built upon ATOMIC and consists of event conceptualization data and abstract knowledge triples. The event conceptualizations are based on head events in ATOMIC, identified through syntactic parsing and matching with rules to search for concept candidates in Probase Wu et al. (2012) and WordNet Miller (1995). The abstract knowledge triples connect conceptualized head events with their non-abstract counterparts from ATOMIC, forming commonsense knowledge at the concept level. Human annotations are used to verify the correctness of some conceptualizations and their resulting abstract commonsense triples. In total, 131K conceptualizations of 7K (45%) ATOMIC head events and 81K (1.3%) conceptualized triples are manually annotated, with a large number remaining unlabeled. Detailed statistics are shown in Table 7. In this paper, we evaluate all models using the test set from the annotated subset as the evaluation data. Meanwhile, we obtain CANDLE distilled models using the training set from the annotated subset to fine-tune discriminative models pre-trained on CANDLE conceptualizations.
B.2 Generative Commonsense Inference
To evaluate COMET, we adopt the same evaluation setting employed by Hwang et al. (2021) for assessing commonsense generative models on the ATOMIC dataset’s test set. We use the entire test set, consisting of 34,689 triples across 23 different commonsense relations, to ensure the robustness of the evaluation. Additionally, we use the full training set to fine-tune models that were pretrained on various CSKBs and CANDLE instantiations.
@@ B.2 Generative Commonsense Inference
B.3 Zero-shot Commonsense QA
@@
Appendix C Implementation Details
This section provides additional information about implementation details in downstream task evaluations.
First, we use the Huggingface Library Wolf et al. (2020) to build all models. We reproduce all baselines according to implementation details described in their original papers. The reported results are consistent with their original papers if the same experiment is included. For CANDLE distilled models, please refer to the subsections (Appendix C.1, C.2, C.3) below.
For methods involving LLMs, we use their instruction fine-tuned versions as the backbone for the baselines. For LLAMA2, the accessed version is meta-llama/Llama-2-7b/13b-chat-hf. For Mistral, we use mistralai/Mistral -7B-Instruct-v0.1. This remains consistent whether we prompt them directly or fine-tune them for downstream tasks, as we have observed that the instruction-finetuned versions generally result in better performance. For ChatGPT, we access it through Microsoft Azure APIs333https://azure.microsoft.com/en-us/products/ai-services/. The code of the accessed version for ChatGPT is gpt-35-turbo-20230515, and for GPT4 is gpt-4-20230515. The maximum generation length is set to 100 tokens for all tasks. For fine-tuning LLAMA2 and Mistral, we use the open code base of LLaMa-Factory444https://github.com/hiyouga/LLaMA-Factory. Please refer to the subsections below for hyperparameter settings.
@@ C.1 CSKB Conceptualization
For VERA, we follow the exact same implementation555https://github.com/liujch1998/vera as released by Liu et al. (2023). To transform our binary classification subtasks into declarative formats, we begin by converting each piece of data into a declarative sentence using predefined natural language templates. Next, we create a corresponding negative statement by simply incorporating the word “not” into the correct sentence. For instance, a pair of statements is: “PersonX enjoys drinking at the bar. Bar is a social gathering place.” and “PersonX enjoys drinking at the bar. Bar is not a social gathering place.” The accessed backbone model is liujch1998/vera, and all other hyperparameter settings follow the default implementation. The model is trained on CANDLE distillation for one epoch and then fine-tuned on AbstractATOMIC for another.
C.2 Generative Commonsense Inference
@@
| DeBERTa-v3-L (CANDLE Distilled) | CANDLE | 81.2 | 69.9 | 80.3 | 65.9 | 78.3 | 74.9 |
@@
| VERA-T5-xxl (CANDLE Distilled) | CANDLE | 73.8 | 64.7 | 77.6 | 59.4 | 71.3 | 69.4 |
@@
C.3 Zero-shot Commonsense QA
For the task of zero-shot commonsense QA, we adopt the code base provided by Wang et al. (2023a)666https://github.com/HKUST-KnowComp/CAR and Liu et al. (2023)777https://github.com/liujch1998/vera to train two CANDLE distilled models. All hyperparameters are kept unchanged from their original implementations as default settings. The models are trained for two epochs using QA pairs obtained from augmented-ATOMIC, including augmentations from ATOMIC-10X, AbstractATOMIC, and CANDLE instantiations.
@@ C.3 Zero-shot Commonsense QA
Appendix D Annotation Details
@@
This paper utilizes expert annotations to assess the quality of distilled conceptualizations and instantiations, as well as evaluate the generations of different models for the COMET downstream task. Four graduate students with ample experience in natural language processing research and expertise in commonsense reasoning are recruited as expert annotators to carry out the annotations. Their participation in the annotation process is voluntary and unpaid, in accordance with local laws, and is considered a contribution to this paper. Detailed instructions are provided to the annotators for each task, ensuring that they understand the requirements thoroughly. For each question, we also highlight the part to be considered by the annotators for their convenience. The annotation process is conducted independently, without any internal discussions among the annotators regarding the results. For each task, two annotators independently vote for each triple, and only when both annotators provide a positive vote will the triple be considered accepted or plausible. To prevent bias and ensure impartial results for CANDLE, the task input is randomly shuffled during the annotation process. As a result, the expert annotators achieve a pairwise agreement (IAA; Landis and Koch, 1977) of 0.80 and a Fleiss-kappa Fleiss (1971) of 0.61, indicating a remarkably high level of internal agreement.
Appendix E Case Study
We present some examples in Table 11 to show conceptualizations and instantiations generated by CANDLE, along with their corresponding critic values assigned by our filtering discriminators. Most of the generations are of high quality, as expected, and can introduce novel events compared to the original knowledge.

### v2

@@ Acknowledgements
We thank the anonymous reviewers and the area chair for their constructive comments. The authors of this paper were supported by the NSFC Fund (U20B2053) from the NSFC of China, the RIF (R6020-19 and R6021-20), and the GRF (16211520 and 16205322) from RGC of Hong Kong. We also thank the support from the UGC Research Matching Grants (RMGS20EG01-D, RMGS20CR11, RMGS20CR12, RMGS20EG19, RMGS20EG21, RMGS23CR05, RMGS23EG08).
@@ Appendix A Distillation Details
These prompts are consistent with our descriptions in Section 4.1 and Section 4.2, where the task description is first presented, followed by human-authored examples, and finally, the event we want to conceptualize or instantiate. We also leverage several tricks in the prompt, such as numbering the examples, generating concepts instead of hypernyms, and keeping the generated responses concise. Finally, we parse the generations via manually defined rules and compile them into a dataset.
Additionally, we introduce some generation settings when prompting LLMs. For ChatGPT, we access it through the official OpenAI APIs222https://chat.openai.com/. The code of the accessed version is gpt-3.5-turbo-0613. We set the temperature to 1.0 and the maximum length for generated tokens to 200. To conceptualize all events in ATOMIC into 20 conceptualizations each, the time required for the distillation process is approximately ten days and the financial budget is around 1500 USD.
For LLAMA2, we access it via the Huggingface Library Wolf et al. (2020). The code of the accessed model is meta-llama/Llama-2-13b-chat-hf333https://huggingface.co/meta-llama/Llama-2-13b-chat-hf. When prompting, we use the Top-k sampling decoding strategy and set . We set the maximum length of generated tokens to 200. The models are hosted on sixteen NVIDIA-V100 GPUs, and the time required to distill the entire dataset is approximately one month.
@@
Similarly, we evaluate the uniqueness of the newly introduced head events resulting from our chain of conceptualization and instantiation. To determine uniqueness, we define an instantiated head event, referred to as , as unique if , where represents the original head event in ATOMIC. The threshold of 0.5 is an empirical threshold. Our empirical results demonstrate that 78.6% of the instantiated events are unique compared to ATOMIC, highlighting the effectiveness of CANDLE in enhancing the semantic coverage of the CSKB.
Appendix B Task Setups
B.1 CSKB Conceptualization
We follow the task definition of He et al. (2024) and Wang et al. (2023b) to formulate the CSKB conceptualization task. Specifically, conceptualizing an event-centric CSKB to derive abstract commonsense knowledge comprises two steps He et al. (2024): event conceptualization and triple conceptualization, which correspond to two subtasks studied in this paper. Denote the triples in the original CSKB as , where , , and are the set of heads, relations, and tails in the original CSKB. The first step only operates on head events without considering the context in and . The goal of event conceptualization is to produce a conceptualized head event from the original head to represent an abstraction of . In the second step, the task is to verify whether the conceptualized head still makes sense in the context of and , as and will further restrict the level of abstractness in . Plausible triples will be considered as valid abstract commonsense knowledge. By enhancing the performance of discriminative models on these tasks, they can function as more precise critic filters and automate the conceptualization process of a CSKB when linked to concept taxonomies.
B.2 Generative Commonsense Inference
The task of generative commonsense inference was studied by both Bosselut et al. (2019) and Hwang et al. (2021). It requires a generative model to complete the tail of a commonsense assertion based on a given pair of head and commonsense relation . In this paper, we follow Hwang et al. (2021) and use ATOMIC as the evaluation benchmark, in which the full testing set is used for model evaluation. The task of COMET is important in the domain of commonsense as it serves as a fundamental component for numerous high-level applications that necessitate commonsense reasoning, such as zero-shot commonsense question answering with self-talk Shwartz et al. (2020) and dynamic graph construction Bosselut et al. (2021), narrative reasoning Peng et al. (2022b), and dialogue generation Tu et al. (2022). Improving COMET can potentially benefit other domains that require commonsense understanding.
B.3 Zero-shot Commonsense QA
The task of zero-shot commonsense QA evaluates a model’s reasoning generalizability on unseen QA entries without any supervision signals from the corresponding annotated training data. Several methods have been proposed to tackle this task, including those by Shwartz et al. (2020); Bosselut et al. (2021); Kim et al. (2022); Shi et al. (2023). The most effective pipeline, as suggested by Ma et al. (2021), injects commonsense knowledge into language models via fine-tuning on QA pairs synthesized from knowledge in CSKBs. During the fine-tuning process, the head and relation of a triple from a CSKB are transformed into a question using natural language prompts, with the tail serving as the correct answer option. Distractors or negative examples are generated by randomly sampling tails from triples that do not share common keywords with the head. This fine-tuning procedure enhances the model’s knowledge not only for QA benchmarks constructed from CSKBs but also improves its ability to answer unseen commonsense questions in a more generalized manner. In this paper, we follow the task definition, model training, and model evaluation pipeline by Ma et al. (2021) to study the impact of distilling student models from CANDLE instantiations. For baselines, we compare models trained on QA pairs synthesized from ATOMIC, ATOMIC-10X, and AbstractATOMIC. For ATOMIC-10X, 0.9 is used as the critic filtering threshold.
@@ B.3 Zero-shot Commonsense QA
Appendix C Dataset Descriptions
This section covers additional details and statistics of datasets and benchmarks used in downstream task evaluations.
C.1 CSKB Conceptualization
In CSKB Conceptualization tasks, we use the AbstractATOMIC He et al. (2024) dataset as the evaluation benchmark. It is a benchmark dataset built upon ATOMIC and consists of event conceptualization data and abstract knowledge triples. The event conceptualizations are based on head events in ATOMIC, identified through syntactic parsing and matching with rules to search for concept candidates in Probase Wu et al. (2012) and WordNet Miller (1995). The abstract knowledge triples connect conceptualized head events with their non-abstract counterparts from ATOMIC, forming commonsense knowledge at the concept level. Human annotations are used to verify the correctness of some conceptualizations and their resulting abstract commonsense triples. In total, 131K conceptualizations of 7K (45%) ATOMIC head events and 81K (1.3%) conceptualized triples are manually annotated, with a large number remaining unlabeled. The data is partitioned by following ATOMIC’s original split of head events. Detailed statistics are shown in Table 9. In this paper, we evaluate all models using the test set from the annotated subset as the evaluation data. Meanwhile, we obtain CANDLE distilled models using the training set from the annotated subset to fine-tune discriminative models pre-trained on CANDLE conceptualizations. Supervised baselines are trained on the training set of AbstractATOMIC, while semi-supervised baselines also leverage the unlabeled data.
C.2 Generative Commonsense Inference
To evaluate COMET, we adopt the same evaluation setting employed by Hwang et al. (2021) for assessing commonsense generative models on the ATOMIC dataset’s test set. We use the entire test set, consisting of 34,689 triples across 23 different commonsense relations, to ensure the robustness of the evaluation. Additionally, we use the full training set to fine-tune models that were pre-trained on various CSKBs and CANDLE instantiations to fit them into the benchmark.
@@ C.2 Generative Commonsense Inference
C.3 Zero-shot Commonsense QA
@@
Appendix D Implementation Details
This section provides additional implementation details in downstream task evaluations.
First, we use the Huggingface444https://huggingface.co/ Library Wolf et al. (2020) to build all models. We reproduce all baselines according to implementation details described in their original papers. The reported results are consistent with their original papers if the same experiment is included. For CANDLE distilled models, please refer to the subsections (Appendix D.1, D.2, D.3) below.
For methods involving LLMs, we use their instruction fine-tuned versions as the backbone for the baselines. For LLAMA2, the accessed version is meta-llama/Llama-2-7b/13b-chat-hf. For Mistral, we use mistralai/Mistral -7B-Instruct-v0.1. This remains consistent whether we prompt them directly or fine-tune them for downstream tasks, as we have observed that the instruction-finetuned versions generally result in better performance. For ChatGPT, we access it through Microsoft Azure APIs555https://azure.microsoft.com/en-us/products/ai-services/. The code of the accessed version for ChatGPT is gpt-35-turbo-20230515, and for GPT4 is gpt-4-20230515. The maximum generation length is set to 100 tokens for all tasks. For fine-tuning LLAMA2 and Mistral, we use the open code base of LLaMa-Factory666https://github.com/hiyouga/LLaMA-Factory. Please refer to the subsections below for hyperparameter settings.
@@ D.1 CSKB Conceptualization
For VERA, we follow the exact same implementation777https://github.com/liujch1998/vera as released by Liu et al. (2023). To transform our binary classification subtasks into declarative formats, we begin by converting each piece of data into a declarative sentence using predefined natural language templates. Next, we create a corresponding negative statement by simply incorporating the word “not” into the correct sentence. For instance, a pair of statements is: “PersonX enjoys drinking at the bar. Bar is a social gathering place.” and “PersonX enjoys drinking at the bar. Bar is not a social gathering place.” The accessed backbone model is liujch1998/vera, and all other hyperparameter settings follow the default implementation. The model is trained on CANDLE distillation for one epoch and then fine-tuned on AbstractATOMIC for another.
D.2 Generative Commonsense Inference
@@
| DeBERTa-v3-L (CANDLE Distilled) | CANDLE | 81.2↑1.6 | 69.9↑0.6 | 80.3↑1.7 | 65.9↑1.9 | 78.3↑0.1 | 74.9↑1.0 |
@@
| VERA-T5-xxl (CANDLE Distilled) | CANDLE | 73.8↑0.6 | 64.7↑1.7 | 77.6↑0.4 | 59.4↑1.2 | 71.3↑3.2 | 69.4↑1.4 |
@@
D.3 Zero-shot Commonsense QA
For the task of zero-shot commonsense QA, we adopt the code base provided by Wang et al. (2023a)888https://github.com/HKUST-KnowComp/CAR and Liu et al. (2023)999https://github.com/liujch1998/vera to train two CANDLE distilled models. All hyperparameters and optimization strategies are kept unchanged from their original implementations as default settings. The models are trained for two epochs using QA pairs obtained from augmented-ATOMIC, including augmentations from ATOMIC-10X, AbstractATOMIC, and CANDLE instantiations.
@@ D.3 Zero-shot Commonsense QA
Appendix E Annotation Details
@@
This paper utilizes expert annotations to assess the quality of distilled conceptualizations and instantiations, as well as evaluate the generations of different models for the COMET downstream task. We follow the annotation setting of Yu et al. (2023); Lu et al. (2024); Cheng et al. (2023), where four graduate students with ample experience in natural language processing research and expertise in commonsense reasoning are recruited as expert annotators to carry out the annotations. Their participation in the annotation process is voluntary and unpaid, in accordance with local laws, and is considered a contribution to this paper. Detailed instructions are provided to the annotators for each task, ensuring that they understand the requirements thoroughly. For CANDLE distillation evaluation, the annotators are asked to determine (1) the correctness of the conceptualizations and instantiations and (2) the plausibility of their formed triples. For COMET generation evaluation, they are asked to determine the plausibility of the generated triples. For each question, we also highlight the part to be considered by the annotators for their convenience. The annotation process is conducted independently, without any internal discussions among the annotators regarding the results. For each task, two annotators independently vote for each triple, and only when both annotators provide a positive vote will the triple be considered accepted or plausible. To prevent bias and ensure impartial results for CANDLE, the task input is randomly shuffled during the annotation process. As a result, the expert annotators achieve a pairwise agreement (IAA; Landis and Koch, 1977) of 0.80 and a Fleiss-kappa Fleiss (1971) of 0.61, indicating a remarkably high level of internal agreement.
Appendix F Ablation Study
In this section, we examine the impact of our critic filters on the ablation of CAN-DLE. Specifically, we investigate the effect of different levels of critic threshold or completely abandoning critic filtering on downstream tasks. We conduct four experiments with different settings, denoted as , where = 0 corresponds to abandoning critic filtering and using all distilled knowledge as complementary training data. For detailed statistics, please refer to Table 1. For each value of , we select the distilled knowledge with a critic score higher than and utilize it as complementary training data to train student models for the three downstream tasks. We employ the same training strategies described in the main body of the paper. In the case of CSKB conceptualization and zero-shot commonsense QA tasks, we utilize DeBERTa-v3-large as the backbone model, with accuracy as the evaluation metric. For COMET, we use GPT2 and evaluate using the BERTScore as the evaluation metric. The results are visualized in Figure 4. Our analysis reveals a consistent trend where higher threshold values yield improved performance, indicating the reliability of our critic filter. However, it is worth noting that setting the threshold above 0.9 may potentially lead to even better performance. Nevertheless, such a trade-off comes with a downside: it reduces the amount of usable knowledge in each distillation round, which can impede the iterative process. The reason for this is that when the number of distilled conceptualizations and instantiations decreases significantly in each round, CANDLE is unable to incorporate new instantiated data for future distillation iterations. As a result, the “convergence” of those high-critic data occurs prematurely in CANDLE.
Appendix G Case Study
We present some examples in Table 13 to show conceptualizations and instantiations generated by CANDLE, along with their corresponding critic values assigned by our critic-filtering discriminators. It can be observed that both ChatGPT and LLAMA2 exhibit the ability to generate high-quality knowledge based on given instructions. Furthermore, they can introduce novel conceptualizations and events during the distillation chain, effectively meeting our expectations of CANDLE. Future works can investigate the feasibility of incorporating conceptualization and abstract knowledge into more downstream tasks, such as complex reasoning Fang et al. (2024); Zhang et al. (2024); Bai et al. (2023) and commonsense knowledge graph denoising Deng et al. (2023).